OpenStudy (anonymous):

Multiply and simplify: quantity 9 + 2 radical 3 times quantity 9 minus 2 radical 3

OpenStudy (campbell_st):

get the binomial pairs \[(9 + 2\sqrt{3})(9 - 2\sqrt{3})\] this expands to the difference of 2 squares (a +b)(a -b) = a^2 - b^2 use this method should get 69 as the answer

OpenStudy (anonymous):

thanks your a real help :)